Maxwell Mining Company's ore reserves are being depleted, so its sales are falling. Also, because its pit is getting deeper each year, its costs are rising. As a result, the company's earnings and dividends are declining at the constant rate of 10% per year. If D0 = $5 and rs = 9%, what is the value of Maxwell Mining's stock? Round your answer to two decimal places.
Answer #1

Current price=D1/(Required return-Growth rate)

=(5(1-0.1))/(0.09-(0.1)

=(4.5/0.19)

=$23.68